OpenStudy (anonymous):

An object experiences an acceleration of -6.8 m/s. as a result it accelerates from 54 m to a complete stop. How much distance did it travel ?

OpenStudy (anonymous):

Let's put out the known variables first

OpenStudy (anonymous):

We know that: Acceleration = -6.8 m/s² V1 = 54 m V2 = 0 D = ?

OpenStudy (anonymous):

You should be able to solve for Distance after rearranging the equationa bit

OpenStudy (anonymous):

V2² = V1² + 2ad V2² - V1² / 2a = d 0 - 54² / 2(-6.8) = d ?

OpenStudy (anonymous):

Omgosh thank you so much . Do you mind helping me with5 mor questions pleeeeeease ?!!

OpenStudy (anonymous):

Ps- you really are the boss ^___^

OpenStudy (anonymous):

jus post them on this website im sure myself and others would love to help